sally bought three chocolate bars and a pack of gum and paid 41.75. jAKE BOUGHT TWO chocolate bars and four packs of gum and paid $2.00. Find the cost of a chocolate bar and the cost of a pack of gum. BRAIN FRIED
Answer by ankor@dixie-net.com(12692)

The first statement has an obvious mistake. Changed 41.75 to 1.75

Let x = price of choc bars, and y = price of gum
:
Write an equation representing each statement, we will use elimination:
:
"sally bought three chocolate bars and a pack of gum and paid 1.75."
3x + 1y = 1.75
:
"jAKE BOUGHT TWO chocolate bars and four packs of gum and paid $2.00."
2x + 4y = 2.00
:
Find the cost of a chocolate bar and the cost of a pack of gum.
:
Multiply the 1st equation by 4 and subtract the 2nd equation from it:
12x + 4y = 7.00
2x + 4y = 2.00
------------------- subtracting eliminates y, find x.
10x + 0y = 5.00
x = 5.00/10
x = .50 for a chocolate bar
:
Find y using the 1st equation, substitute .50 for x
3(.50) + 1y = 1.75
1.50 + 1y = 1.75
y = 1.75 - 1.50
y = .25 for gum
:
Check our solutions in the 2nd equation:
2(.50) + 4(.25) =
1.00 + 1.00 = 2.00
:
How about this, has your brain recovered somewhat? Any questions?